# FreDN: Spectral Disentanglement for Time Series Forecasting via Learnable Frequency Decomposition

### v1

@@ 14. $\hat{Y}=\operatorname{RevIN}^{-1}(\hat{Y}_{\text{season}}+\hat{Y}_{\text{trend}})$
Following the protocol proposed in iTransformer (liu2024itransformer), we split all datasets into training, validation, and testing subsets strictly based on chronological order to avoid information leakage. The ETT datasets are divided in a 6:2:2 ratio, while all other datasets follow a 7:1:2 split. We set the forecasting length to $\tau\in\{96,192,336,720\}$ for all models and datasets to ensure comparability. Instead of using a fixed look-back window size, we vary the historical window length $L\in\{96,192,336,512,720\}$, and report the best performance achieved over different values of $L$ for each forecasting horizon. This design allows us to better evaluate the ability of models to utilize long-range temporal information.
@@
All experiments are implemented in PyTorch (paszke2019pytorch) and conducted on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. The average training time is shown in Figure 9 The ADAM optimizer (kingma2014adam) is used across all experiments. We use a fixed random seed (e.g., 2025) for reproducibility unless otherwise specified.
For our proposed FreDN model, we employ the frequency-domain MAE loss introduced in FreDF (wang2025fredf), which is defined as
@@
Reversible Instance Normalization (RevIN) (kim2021reversible) is a commonly adopted technique to reduce distributional discrepancies between training and testing data. It removes per-instance temporal trends before modeling and restores them after prediction. Following prior works (liu2024itransformer; wang2024timemixer), we adopt RevIN in our model. The procedure involves three steps:
@@
MovDN follows the moving average decomposition strategy proposed in Autoformer (wu2021autoformer), where the trend component is extracted via a symmetric padding and average pooling operation. Formally, given an input sequence $X\in\mathbb{R}^{C\times L\times d}$, the trend is obtained by applying a 1D average pooling over the temporal dimension with a fixed kernel size (we choose 25 following (wang2024timemixer)), and the seasonal component is computed as the residual between the original input and the smoothed trend. The implementation matches that of series_decomp in the original Autoformer repository.
@@
Trabelsi et al. (trabelsi2018deep) propose that activation and dropout can be applied separately to the real and imaginary parts. For LayerNorm, two design options are presented: (1) treating the complex tensor as a joint distribution and whitening it using the full covariance matrix between the real and imaginary parts; and (2) applying standard LayerNorm to the real and imaginary parts independently.
To further evaluate the utility of our proposed ReIm Block, we introduce Decorr-CL and Separate-CL as enhanced variants of the Complex-Linear structure, which implement complex-valued LayerNorm, activation functions, and dropout mechanisms following the framework established by Trabelsi et al. (trabelsi2018deep) [^5] https://github.com/ChihebTrabelsi/deep˙complex˙networks. We present the comparative results in  Appendix C.
@@
In both cases, we follow the convention of Trabelsi et al. (trabelsi2018deep) where GELU activation and dropout are applied separately to the real and imaginary parts. The implementations used in our experiments mirrors our ResMLP exactly in topology, with Linear, LayerNorm, GELU, and dropout replaced by their complex counterparts as described above.
Appendix Appendix C Full Results
@@
Table 5: Multivariate forecasting results with horizon $H\in\{96,192,336,720\}$ for all datasets. We evaluate each model under multiple lookback window lengths $\{96,192,336,512,720\}$ and report the best performance among them. The best results are highlighted in bold red, and the second best are underlined blue.
@@
Models FreDN SOFTS FreDF iTransformer TimeMixer FreTS PatchTST DLinear
(Ours) (2024) (2025) (2024) (2023) (2023) (2023) (2023)
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.355 0.387 0.382 0.401 0.383 0.405 0.387 0.405 0.386 0.401 0.395 0.410 0.378 0.409 0.367 0.395
192 0.391 0.410 0.414 0.429 0.476 0.466 0.427 0.436 0.421 0.430 0.436 0.442 0.417 0.430 0.400 0.417
336 0.412 0.426 0.438 0.443 0.517 0.492 0.449 0.459 0.439 0.453 0.476 0.467 0.438 0.445 0.430 0.438
720 0.429 0.457 0.452 0.470 0.552 0.529 0.492 0.492 0.476 0.475 0.547 0.532 0.464 0.475 0.468 0.489
Avg 0.397 0.420 0.421 0.436 0.482 0.473 0.438 0.448 0.430 0.440 0.463 0.463 0.424 0.440 0.416 0.435
ETTh2
96 0.266 0.328 0.285 0.348 0.293 0.343 0.299 0.350 0.280 0.343 0.312 0.373 0.282 0.341 0.290 0.357
192 0.320 0.368 0.355 0.393 0.364 0.398 0.377 0.399 0.351 0.389 0.392 0.428 0.350 0.385 0.399 0.430
336 0.352 0.396 0.388 0.418 0.396 0.429 0.415 0.432 0.380 0.419 0.491 0.489 0.373 0.404 0.480 0.481
720 0.383 0.426 0.415 0.445 0.429 0.450 0.416 0.443 0.425 0.445 0.653 0.577 0.398 0.432 0.762 0.620
Avg 0.330 0.380 0.361 0.401 0.371 0.405 0.377 0.406 0.359 0.399 0.462 0.467 0.351 0.391 0.483 0.472
ETTm1
96 0.284 0.334 0.303 0.353 0.303 0.348 0.302 0.353 0.300 0.347 0.307 0.351 0.304 0.350 0.300 0.343
192 0.319 0.361 0.336 0.376 0.338 0.369 0.345 0.378 0.340 0.374 0.348 0.377 0.349 0.377 0.335 0.365
336 0.354 0.382 0.371 0.396 0.367 0.390 0.379 0.403 0.374 0.397 0.383 0.399 0.378 0.392 0.370 0.388
720 0.402 0.412 0.428 0.430 0.492 0.474 0.439 0.437 0.435 0.425 0.444 0.437 0.422 0.418 0.415 0.416
Avg 0.340 0.372 0.360 0.389 0.375 0.395 0.366 0.393 0.362 0.386 0.370 0.391 0.363 0.384 0.355 0.378
ETTm2
96 0.156 0.244 0.175 0.258 0.164 0.254 0.175 0.266 0.171 0.257 0.172 0.261 0.165 0.256 0.166 0.250
192 0.211 0.283 0.244 0.306 0.226 0.296 0.238 0.312 0.227 0.298 0.242 0.311 0.225 0.297 0.221 0.291
336 0.264 0.318 0.295 0.341 0.277 0.333 0.289 0.341 0.279 0.331 0.308 0.356 0.274 0.328 0.273 0.328
720 0.341 0.370 0.379 0.396 0.369 0.388 0.377 0.397 0.364 0.383 0.373 0.404 0.354 0.383 0.362 0.393
Avg 0.243 0.304 0.273 0.325 0.259 0.318 0.270 0.329 0.260 0.317 0.274 0.333 0.254 0.316 0.255 0.316
Weather
96 0.139 0.186 0.162 0.213 0.156 0.206 0.163 0.213 0.150 0.204 0.151 0.210 0.159 0.211 0.168 0.228
192 0.183 0.230 0.207 0.254 0.204 0.250 0.203 0.249 0.195 0.244 0.195 0.253 0.203 0.251 0.211 0.266
336 0.232 0.270 0.259 0.293 0.251 0.287 0.255 0.290 0.245 0.281 0.242 0.293 0.258 0.293 0.259 0.310
720 0.305 0.324 0.338 0.345 0.322 0.337 0.326 0.338 0.323 0.335 0.309 0.340 0.320 0.339 0.314 0.351
Avg 0.215 0.253 0.242 0.276 0.233 0.270 0.237 0.272 0.228 0.266 0.224 0.274 0.235 0.274 0.238 0.289
Traffic
96 0.126 0.219 0.130 0.226 0.130 0.226 0.131 0.227 0.133 0.226 0.135 0.275 0.133 0.270 0.137 0.286
192 0.144 0.236 0.149 0.245 0.147 0.241 0.153 0.249 0.154 0.245 0.151 0.288 0.149 0.278 0.151 0.289
336 0.159 0.252 0.166 0.262 0.163 0.257 0.169 0.265 0.169 0.264 0.168 0.304 0.168 0.283 0.166 0.296
720 0.191 0.283 0.197 0.293 0.199 0.290 0.191 0.285 0.206 0.296 0.208 0.334 0.207 0.303 0.201 0.318
Avg 0.155 0.247 0.161 0.256 0.160 0.253 0.161 0.257 0.165 0.258 0.166 0.300 0.164 0.283 0.164 0.297
Traffic
96 0.353 0.247 0.348 0.243 0.375 0.269 0.351 0.257 0.359 0.251 0.391 0.275 0.374 0.270 0.399 0.286
192 0.370 0.255 0.362 0.255 0.395 0.281 0.365 0.265 0.371 0.261 0.412 0.288 0.389 0.278 0.408 0.289
336 0.381 0.259 0.378 0.263 0.418 0.296 0.384 0.273 0.388 0.265 0.435 0.304 0.404 0.283 0.422 0.296
720 0.420 0.283 0.414 0.279 0.456 0.316 0.415 0.287 0.427 0.291 0.488 0.334 0.442 0.303 0.462 0.318
Avg 0.381 0.261 0.376 0.260 0.411 0.290 0.379 0.270 0.386 0.267 0.431 0.300 0.402 0.283 0.423 0.297
1st Count 30 32 5 4 0 0 1 0 0 0 0 0 0 0 0 0
2nd Count 1 4 3 3 6 4 4 1 3 6 3 0 6 7 10 13
@@
Comparison Metric
Wilcoxon
Statistic
p-value Significant
@@
FreDN (Ours) MSE 0.397 0.330 0.340 0.243 0.215 0.155 0.381
MAE 0.420 0.380 0.372 0.304 0.253 0.247 0.261
TimeMixer++ (2025) MSE 0.419 0.339 0.369 0.269 0.226 0.165 0.416
MAE 0.432 0.380 0.378 0.320 0.262 0.253 0.264
@@
In Table 8, we report the full comparison results for the two enhanced Complex-Linear structures introduced in  Appendix B: Decorr-CL and Separate-CL. These two structures extend the base Complex-Linear structure by incorporating complex-valued LayerNorm, activation functions, and dropout, following the formulation proposed by Trabelsi et al. (trabelsi2018deep). Specifically, Decorr-CL computes the full second-order statistics of the real and imaginary components, normalizes using their joint covariance structure, and applies separate trainable affine parameters. This makes Decorr-CL heavier in terms of parameter count compared to the base Complex-Linear.
@@
Table 8: Performance comparison between ReIm Block, Complex-Linear, Decorr-CL, and Separate-CL. The reported results are averaged over four prediction lengths $\tau\in\{96,192,336,720\}$.
@@
Proof of Theorem: Expressiveness of ReIm Block
Theorem (Expressiveness of ReIm Block). Let $\tilde{X}=[\tilde{X}_{1},\dots,\tilde{X}_{d}]^{\top}\in\mathbb{C}^{d}$ be a complex-valued input vector.
@@
As shown in Hoffman and Kunze (hoffman1971linear), two vectors $\tilde{X}_{i},\tilde{X}_{j}\in\mathbb{C}$ are linearly independent over $\mathbb{R}$ if and only if their ratio is not real, i.e.,

### v2

@@ 14. $\hat{Y}=\operatorname{RevIN}^{-1}(\hat{Y}_{\text{season}}+\hat{Y}_{\text{trend}})$
Following the protocol proposed in iTransformer (Liu et al. 2024), we split all datasets into training, validation, and testing subsets strictly based on chronological order to avoid information leakage. The ETT datasets are divided in a 6:2:2 ratio, while all other datasets follow a 7:1:2 split. We set the forecasting length to $\tau\in\{96,192,336,720\}$ for all models and datasets to ensure comparability. Instead of using a fixed look-back window size, we vary the historical window length $L\in\{96,192,336,512,720\}$, and report the best performance achieved over different values of $L$ for each forecasting horizon. This design allows us to better evaluate the ability of models to utilize long-range temporal information.
@@
All experiments are implemented in PyTorch (Paszke et al. 2019) and conducted on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. The average training time is shown in Figure 9 The ADAM optimizer (Kingma and Ba 2015) is used across all experiments. We use a fixed random seed (e.g., 2025) for reproducibility unless otherwise specified.
For our proposed FreDN model, we employ the frequency-domain MAE loss introduced in FreDF (Wang et al. 2025a), which is defined as
@@
Reversible Instance Normalization (RevIN) (Kim et al. 2021) is a commonly adopted technique to reduce distributional discrepancies between training and testing data. It removes per-instance temporal trends before modeling and restores them after prediction. Following prior works (Liu et al. 2024; Wang et al. 2024), we adopt RevIN in our model. The procedure involves three steps:
@@
MovDN follows the moving average decomposition strategy proposed in Autoformer (Wu et al. 2021), where the trend component is extracted via a symmetric padding and average pooling operation. Formally, given an input sequence $X\in\mathbb{R}^{C\times L\times d}$, the trend is obtained by applying a 1D average pooling over the temporal dimension with a fixed kernel size (we choose 25 following (Wang et al. 2024)), and the seasonal component is computed as the residual between the original input and the smoothed trend. The implementation matches that of series_decomp in the original Autoformer repository.
@@
Trabelsi et al. (Trabelsi et al. 2018) propose that activation and dropout can be applied separately to the real and imaginary parts. For LayerNorm, two design options are presented: (1) treating the complex tensor as a joint distribution and whitening it using the full covariance matrix between the real and imaginary parts; and (2) applying standard LayerNorm to the real and imaginary parts independently.
To further evaluate the utility of our proposed ReIm Block, we introduce Decorr-CL and Separate-CL as enhanced variants of the Complex-Linear structure, which implement complex-valued LayerNorm, activation functions, and dropout mechanisms following the framework established by Trabelsi et al. (Trabelsi et al. 2018) [^5] https://github.com/ChihebTrabelsi/deep˙complex˙networks. We present the comparative results in Appendix C,  C.
@@
In both cases, we follow the convention of Trabelsi et al. (Trabelsi et al. 2018) where GELU activation and dropout are applied separately to the real and imaginary parts. The implementations used in our experiments mirrors our ResMLP exactly in topology, with Linear, LayerNorm, GELU, and dropout replaced by their complex counterparts as described above.
Appendix C Full Results
@@
Table 5: Full results of multivariate forecasting with horizon $H\in\{96,192,336,720\}$ for all datasets. We evaluate each model under multiple lookback window lengths $\{96,192,336,512,720\}$ and report the best performance. The best results are highlighted in bold red, and the second best are underlined blue.
@@
Models FreDN SOFTS FreDF iTransformer TimeMixer FreTS PatchTST DLinear (Ours) (2024) (2025) (2024) (2023) (2023) (2023) (2023) Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 96 0.355 0.387 0.383 0.405 0.382 0.401 0.387 0.405 0.386 0.401 0.395 0.410 0.378 0.409 0.367 0.395 192 0.391 0.410 0.476 0.466 0.414 0.429 0.427 0.436 0.421 0.430 0.436 0.442 0.417 0.430 0.400 0.417 336 0.412 0.426 0.517 0.492 0.438 0.443 0.449 0.459 0.439 0.453 0.476 0.467 0.438 0.445 0.430 0.438 720 0.429 0.457 0.552 0.529 0.452 0.470 0.492 0.492 0.476 0.475 0.547 0.532 0.464 0.475 0.468 0.489 Avg 0.397 0.420 0.482 0.473 0.421 0.436 0.438 0.448 0.430 0.440 0.463 0.463 0.424 0.440 0.416 0.435 ETTh2 96 0.266 0.328 0.293 0.343 0.285 0.348 0.299 0.350 0.280 0.343 0.312 0.373 0.282 0.341 0.290 0.357 192 0.320 0.368 0.364 0.398 0.355 0.393 0.377 0.399 0.351 0.389 0.392 0.428 0.350 0.385 0.399 0.430 336 0.352 0.396 0.396 0.429 0.388 0.418 0.415 0.432 0.380 0.419 0.491 0.489 0.373 0.404 0.480 0.481 720 0.383 0.426 0.429 0.450 0.415 0.445 0.416 0.443 0.425 0.445 0.653 0.577 0.398 0.432 0.762 0.620 Avg 0.330 0.380 0.371 0.405 0.361 0.401 0.377 0.406 0.359 0.399 0.462 0.467 0.351 0.391 0.483 0.472 ETTm1 96 0.284 0.334 0.303 0.348 0.303 0.353 0.302 0.353 0.300 0.347 0.307 0.351 0.304 0.350 0.300 0.343 192 0.319 0.361 0.338 0.369 0.336 0.376 0.345 0.378 0.340 0.374 0.348 0.377 0.349 0.377 0.335 0.365 336 0.354 0.382 0.367 0.390 0.371 0.396 0.379 0.403 0.374 0.397 0.383 0.399 0.378 0.392 0.370 0.388 720 0.402 0.412 0.492 0.474 0.428 0.430 0.439 0.437 0.435 0.425 0.444 0.437 0.422 0.418 0.415 0.416 Avg 0.340 0.372 0.375 0.395 0.360 0.389 0.366 0.393 0.362 0.386 0.370 0.391 0.363 0.384 0.355 0.378 ETTm2 96 0.156 0.244 0.164 0.254 0.175 0.258 0.175 0.266 0.171 0.257 0.172 0.261 0.165 0.256 0.166 0.250 192 0.211 0.283 0.226 0.296 0.244 0.306 0.238 0.312 0.227 0.298 0.242 0.311 0.225 0.297 0.221 0.291 336 0.264 0.318 0.277 0.333 0.295 0.341 0.289 0.341 0.279 0.331 0.308 0.356 0.274 0.328 0.273 0.328 720 0.341 0.370 0.369 0.388 0.379 0.396 0.377 0.397 0.364 0.383 0.373 0.404 0.354 0.383 0.362 0.393 Avg 0.243 0.304 0.259 0.318 0.273 0.325 0.270 0.329 0.260 0.317 0.274 0.333 0.254 0.316 0.255 0.316 Weather 96 0.139 0.186 0.156 0.206 0.162 0.213 0.163 0.213 0.150 0.204 0.151 0.210 0.159 0.211 0.168 0.228 192 0.183 0.230 0.204 0.250 0.207 0.254 0.203 0.249 0.195 0.244 0.195 0.253 0.203 0.251 0.211 0.266 336 0.232 0.270 0.251 0.287 0.259 0.293 0.255 0.290 0.245 0.281 0.242 0.293 0.258 0.293 0.259 0.310 720 0.305 0.324 0.322 0.337 0.338 0.345 0.326 0.338 0.323 0.335 0.309 0.340 0.320 0.339 0.314 0.351 Avg 0.215 0.253 0.233 0.270 0.242 0.276 0.237 0.272 0.228 0.266 0.224 0.274 0.235 0.274 0.238 0.289 Electricity 96 0.126 0.219 0.130 0.226 0.130 0.226 0.131 0.227 0.133 0.226 0.135 0.275 0.133 0.270 0.137 0.286 192 0.144 0.236 0.147 0.241 0.149 0.245 0.153 0.249 0.154 0.245 0.151 0.288 0.149 0.278 0.151 0.289 336 0.159 0.252 0.163 0.257 0.166 0.262 0.169 0.265 0.169 0.264 0.168 0.304 0.168 0.283 0.166 0.296 720 0.191 0.283 0.199 0.290 0.197 0.293 0.191 0.285 0.206 0.296 0.208 0.334 0.207 0.303 0.201 0.318 Avg 0.155 0.247 0.160 0.253 0.161 0.256 0.161 0.257 0.165 0.258 0.166 0.300 0.164 0.283 0.164 0.297 Traffic 96 0.353 0.247 0.375 0.269 0.348 0.243 0.351 0.257 0.359 0.251 0.391 0.275 0.374 0.270 0.399 0.286 192 0.370 0.255 0.395 0.281 0.362 0.255 0.365 0.265 0.371 0.261 0.412 0.288 0.389 0.278 0.408 0.289 336 0.381 0.259 0.418 0.296 0.378 0.263 0.384 0.273 0.388 0.265 0.435 0.304 0.404 0.283 0.422 0.296 720 0.420 0.283 0.456 0.316 0.414 0.279 0.415 0.287 0.427 0.291 0.488 0.334 0.442 0.303 0.462 0.318 Avg 0.381 0.261 0.411 0.290 0.376 0.260 0.379 0.270 0.386 0.267 0.431 0.300 0.402 0.283 0.423 0.297 1st Count 30 32 0 0 5 4 1 0 0 0 0 0 0 0 0 0 2nd Count 1 3 6 4 3 3 4 1 3 6 3 0 6 7 10 13
@@
Comparison Metric Wilcoxon Statistic p-value Significant
@@
FreDN MSE 0.397 0.330 0.340 0.243 0.215 0.155 0.381
(Ours) MAE 0.420 0.380 0.372 0.304 0.253 0.247 0.261
TimeMixer++ MSE 0.419 0.339 0.369 0.269 0.226 0.165 0.416
(2025) MAE 0.432 0.380 0.378 0.320 0.262 0.253 0.264
@@
In Table 8, we report the full comparison results for the two enhanced Complex-Linear structures introduced in Appendix B: Decorr-CL and Separate-CL. These two structures extend the base Complex-Linear structure by incorporating complex-valued LayerNorm, activation functions, and dropout, following the formulation proposed by Trabelsi et al. (Trabelsi et al. 2018). Specifically, Decorr-CL computes the full second-order statistics of the real and imaginary components, normalizes using their joint covariance structure, and applies separate trainable affine parameters. This makes Decorr-CL heavier in terms of parameter count compared to the base Complex-Linear.
@@
Table 8: Performance comparison between Dual-Complex, Complex-Linear, Decorr-CL, and Separate-CL. The reported results are averaged over four prediction lengths $\tau\in\{96,192,336,720\}$.
@@ Appendix D Proof
Proof of Theorem: ReIm Block
Theorem (ReIm Block). Let $\tilde{X}=[\tilde{X}_{1},\dots,\tilde{X}_{d}]^{\top}\in\mathbb{C}^{d}$ be a complex-valued input vector.
@@
As shown in Hoffman and Kunze (Hoffmann and Kunze 1971), two vectors $\tilde{X}_{i},\tilde{X}_{j}\in\mathbb{C}$ are linearly independent over $\mathbb{R}$ if and only if their ratio is not real, i.e.,